# MuDPT: Multi-modal Deep-symphysis Prompt Tuning for Large Pre-trained Vision-Language Models

### v1

@@ Abstract
Prompt tuning, like CoOp, has recently shown promising vision recognizing and transfer learning ability on various downstream tasks with the emergence of large pre-trained vision-language models like CLIP. However, we identify that existing uni-modal prompt tuning approaches may result in sub-optimal performance since this uni-modal design breaks the original alignment of textual and visual representations in the pre-trained model. Inspired by the nature of pre-trained vision-language models, we aim to achieve completeness in prompt tuning and propose a novel approach called Multi-modal Deep-symphysis Prompt Tuning, dubbed as MuDPT, which extends independent multi-modal prompt tuning by additionally learning a model-agnostic transformative network to allow deep hierarchical bi-directional prompt fusion. We evaluate the effectiveness of MuDPT on few-shot vision recognition and out-of-domain generalization tasks. Compared with the state-of-the-art methods, MuDPT achieves better recognition and generalization ability with an apparent margin thanks to synergistic alignment of textual and visual representations. Our code is available at: https://github.com/Mechrev0/MuDPT.
@@ I Introduction
Figure 1: Comparison between our approach: MuDPT and existing prompt tuning approaches: text prompt tuning and visual prompt tuning.
With the VL-PTMs demonstrating attractive generalization capability, the community sets out to explore optional solutions for migrating these models to downstream tasks [6, 7, 8, 9], e.g. visual grounding [6], visual entailment [7], visual recognition [8]. An effective approach is to mitigate the gap between the form of pre-training and downstream task by leveraging carefully determined manual prompts (e.g., “a photo of a [CLASS].”) injected into pre-trained models [4]. Nevertheless, manual prompt engineering is an intuitive and experience-oriented work, and usually issues in sub-optimal performance [8]. Some prompt tuning models are developed for automated prompt engineering, such as CoOp [8], CoCoOp [10], VPT [11], etc. However, as shown in Fig. 1, the previous work has focused on uni-modal prompt tuning in textual or visual branch of VL-PTMs, while multi-modal prompt tuning in the two modalities simultaneously remains an unexplored area in the literature.
@@
To achieve completeness in prompt tuning approaches for better adaption of VL-PTMs, we present Multi-modal Deep-symphysis Prompt Tuning, MuDPT, to adequately fine-tune the textual and visual representations by further learning a modality-agnostic transformative network. We evaluate the effectiveness of our approach on few-shot visual recognition and out-of-domain generalization tasks. Compared with current state-of-the-art methods: CoOp [8] and CoCoOp [10], we obtain better few-shot visual recognition performance with an obvious margin. On base-to-new generalization experiment, our method MuDPT achieves an absotute average gain of 11.28% on new classes and 6.7% on harmonic-mean compared with CoOp. Further, MuDPT demonstrates beneficial transferability in cross-dataset evaluation and domain generalization settings, resulting in better trade-off compared with existing approaches. Extensive experiments demonstrates that MuDPT achieves better few-shot visual recognition and generalization performance thanks to the synergistic alignment of textual and visual representations during prompt tuning on downstream datasets.
@@
  • To achieve cross-modality prompt transformation and fusion, We propose a light modality-agnostic transformative network. It builds the synergistic bridge between textual and visual prompts.
@@ II-B2 Deep Text Prompt Tuning
We conduct deep text prompt tuning in the language branch of CLIP, seen in the upper part of Fig. 3. In our approach, the input embeddings is the concatenation of textual prompts in the embedding layer $T_{0}\in\mathbb{R}^{n\times d_{t}}$ and the word embedding of a specific class name, which is reformulated as $[T_{0},W_{0},w^{eos}_{0}]\in\mathbb{R}^{N\times d_{t}}$ where ${W_{0}=[w^{n+1}_{0},...,w^{N-1}_{0}]}$ is the word embedding of the class name and $w^{eos}_{0}$ is the embedding of $[EOS]$ token. Textual prompts and input embeddings are then further injected into each Transformer layer up to $L^{th}$ layer:
@@ III-B Domain Generalization
Since prompt tuning on a specific dataset risks learning detrimental disturbance in text and image representations which may impede the generalization ability on out-of-distribution (OOD) data, as suggested in [10]. In this section, we aim to evaluate the robustness of MuDPT on distribution shift.
@@
Domain generalization. We measure the generalization ability of MuDPT to OOD data by transferring the prompts learned on ImageNet to the other compatible datasets: ImageNet-V2, ImageNet-Sketch, ImageNet-A, and ImageNet-R. We observe that MuDPT obtains better generalization on 3 of 4 datasets as indicated in Table IV. While obtaining the highest accuracy of 72.13 on the source dataset, we achieve accuracy improvements of 0.85/0.56/0.84 on ImageNet-V2, ImageNet-Sketch, ImageNet-R separately. This demonstrates that MuDPT is a robust prompt tuning approach among the existing methods in domain generalization setting thanks to multi-modal deep-symphysis prompt fusion.

### v2

@@ Abstract
Prompt tuning, like CoOp, has recently shown promising vision recognizing and transfer learning ability on various downstream tasks with the emergence of large pre-trained vision-language models like CLIP. However, we identify that existing uni-modal prompt tuning approaches may result in sub-optimal performance since this uni-modal design breaks the original alignment of textual and visual representations in the pre-trained model. Inspired by the nature of pre-trained vision-language models, we aim to achieve completeness in prompt tuning and propose a novel approach called Multi-modal Deep-symphysis Prompt Tuning, dubbed as MuDPT, which extends independent multi-modal prompt tuning by additionally learning a model-agnostic transformative network to allow deep hierarchical bi-directional prompt fusion. We evaluate the effectiveness of MuDPT on few-shot vision recognition and out-of-domain generalization tasks. Compared with the state-of-the-art methods, MuDPT achieves better recognition and generalization ability with an apparent margin thanks to synergistic alignments of textual and visual representations.
@@ I Introduction
Figure 1: Comparison between our approach (MuDPT) and existing prompt tuning approaches (text prompt tuning and visual prompt tuning).
With the VL-PTMs demonstrating attractive generalization capability, the community sets out to explore optional solutions for migrating these models to downstream tasks [6, 7, 8, 9], e.g., visual grounding [6], visual entailment [7], visual recognition [8]. An effective approach is to mitigate the gap between the form of pre-training and downstream task by leveraging carefully determined manual prompts (e.g., “a photo of a [CLASS].”) injected into pre-trained models [4]. Nevertheless, manual prompt engineering is an intuitive and experience-oriented work, and usually issues in sub-optimal performance [8]. Some prompt tuning models are developed for automated prompt engineering, such as CoOp [8], CoCoOp [10], VPT [11], etc. However, as shown in Fig. 1, the previous work has focused on uni-modal prompt tuning in textual or visual branch of VL-PTMs, while multi-modal prompt tuning in the two modalities simultaneously remains an unexplored area in the literature.
@@
To achieve completeness in prompt tuning approaches for better adaption of VL-PTMs, we present Multi-modal Deep-symphysis Prompt Tuning, MuDPT, to adequately fine-tune the textual and visual representations by further learning a modality-agnostic transformative network. We evaluate the effectiveness of our approach on few-shot visual recognition and out-of-domain generalization tasks. Compared with current state-of-the-art methods: CoOp [8] and CoCoOp [10], we obtain better few-shot visual recognition performance with an obvious margin. On base-to-new generalization experiment, our method MuDPT achieves an absolute average gain of 11.28% on new classes and 6.7% on harmonic-mean compared CoOp. Further, MuDPT demonstrates beneficial transferability in cross-dataset evaluation and domain generalization settings, resulting in better trade-off compared with existing approaches. Extensive experiments demonstrates that MuDPT achieves better few-shot visual recognition and generalization performance thanks to the synergistic alignment of textual and visual representations during prompt tuning on downstream datasets.
@@
  • To achieve cross-modality prompt transformation and fusion, We propose a light modality-agnostic transformative network. It builds the synergistic bridge between text and image prompts.
@@ II-B2 Deep Text Prompt Tuning
We conduct deep text prompt tuning in the language branch of CLIP, seen in the upper part of Fig. 3. In our approach, the input embedding is the concatenation of textual prompts in the embedding layer $T_{0}\in\mathbb{R}^{n\times d_{t}}$ and the word embedding of a specific class name, which is reformulated as $[T_{0},W_{0},w^{eos}_{0}]\in\mathbb{R}^{N\times d_{t}}$ where ${W_{0}=[w^{n+1}_{0},...,w^{N-1}_{0}]}$ is the word embedding of the class name and $w^{eos}_{0}$ is the embedding of $[EOS]$ token. Textual prompts and input embedding are then further injected into each Transformer layer up to $L^{th}$ layer:
@@ III-B Domain Generalization
Since prompt tuning on a specific dataset risks learning detrimental disturbance in text and image representations which may impede the generalization ability on out-of-distribution (OOD) data, as suggested in [10]. In this section, we aim to evaluate the robustness of MuDPT to distribution shift.
@@
Domain generalization. We measure the generalization ability of MuDPT to OOD data by transferring the prompts learned on ImageNet to the other compatible datasets: ImageNet-V2, ImageNet-Sketch, ImageNet-A, and ImageNet-R. We observe that MuDPT obtains better generalization on 3 of 4 datasets as indicated in Table IV. While obtaining the highest accuracy of 72.13 on the source dataset, we achieve accuracy improvements of 0.85/0.56/0.84 on ImageNet-V2, ImageNet-Sketch, ImageNet-R separately. This demonstrates that MuDPT is a robust prompt tuning approach among the existing methods in domain generalization setting thanks to multi-modal deep-symphysis prompts fusion.